# SimBEV: A Synthetic Multi-Task Multi-Sensor Driving Data Generation Tool and Dataset

### v1

@@ Abstract
Bird’s-eye view (BEV) perception for autonomous driving has garnered significant attention in recent years, in part because BEV representation facilitates the fusion of multi-sensor data. This enables a variety of perception tasks including BEV segmentation, a concise view of the environment that can be used to plan a vehicle’s trajectory. However, this representation is not fully supported by existing datasets, and creation of new datasets can be a time-consuming endeavor. To address this problem, in this paper we introduce SimBEV, an extensively configurable and scalable randomized synthetic data generation tool that incorporates information from multiple sources to capture accurate BEV ground truth data, supports a comprehensive array of sensors, and enables a variety of perception tasks including BEV segmentation and 3D object detection. We use SimBEV to create the SimBEV dataset, a large collection of annotated perception data from diverse driving scenarios.
@@ 1 Introduction
Autonomous driving promises a future with safer, cleaner, more efficient, and more reliable transportation systems [11, 58]. As the development of autonomous vehicle technology has accelerated in recent years, so has the need for perception algorithms capable of understanding complex driving scenarios in diverse environments [73, 7]. High-quality driving datasets have been at the center of the recent progress, serving as a foundation for training and benchmarking novel perception algorithms. It is essential for such datasets to encompass a large variety of driving scenarios and encapsulate a diverse set of road types, weather conditions, and traffic patterns, so perception models can effectively generalize to real-world situations [32, 53, 15].
@@
As essential in this context is multimodal sensor fusion, which plays an important role in enhancing the performance of perception algorithms by compensating for weaknesses of one modality with the strengths of others [77, 5, 67]. Sensor fusion improves an autonomous vehicle’s understanding of its surrounding environment [59] (especially in adverse weather conditions [1]), enables robust decision-making in dynamic scenarios [19, 50], and opens the door to perception models that can accomplish multiple tasks simultaneously [35, 43, 47, 20]. Consequently, it is imperative for driving datasets to support a wide array of sensors and various computer vision tasks (such as object detection, semantic segmentation, and tracking) to facilitate the development of multi-faceted perception systems that leverage the strength of multiple sensing modalities.
@@
Bird’s-eye view (BEV) perception has garnered significant attention in recent years for two primary reasons [40]. First, BEV representation is conducive to the fusion of information from different modalities, perspectives, and agents, and extracted BEV features can be used for various perception tasks. Second, BEV segmentation offers a concise, geometrically accurate and semantically rich view of the surrounding environment that can be directly used by motion planning, behavior prediction, and control algorithms. These two factors have given rise to perception algorithms that utilize BEV representation for 3D object detection, BEV segmentation, or both [35, 20, 61, 66, 30, 21, 23, 27, 72, 65, 41, 57, 70, 3, 33, 42, 14, 36, 17, 60, 28, 39, 26, 75, 10, 13, 45, 64, 69, 25, 74].
@@
Figure 1: A data sample generated by SimBEV. The left half depicts a 360-degree view of the ego vehicle’s surroundings using different camera types (from top to bottom RGB, sematic segmentation, instance segmentation, depth, and optical flow cameras, respectively). On the right half, views of lidar, semantic lidar, radar, and the BEV ground truth are shown from top to bottom, respectively. Some images also contain 3D object bounding boxes.
@@
However, despite the growing interest in BEV perception, few existing datasets support BEV segmentation. In the ones that do, either the BEV ground truth is limited to static map elements (drivable area, pedestrian crossing, etc.) [2], or BEV ground truth is only provided for a small window around the ego vehicle, obtained through combining map elements with select 3D object bounding boxes [16].
@@
Creating a new dataset to address this gap is a challenging, time-consuming, and resource-intensive endeavor. Real-world driving data has to be collected over a long period of time to ensure weather conditions and traffic patterns present in the dataset are diverse, and requires (at least partly) labor-intensive hand annotation [56, 32]. As for synthetic datasets, they often consist of carefully designed scenarios that may not capture the full diversity of the driving environment, and may not be as challenging for perception algorithms as their creators think they are. Aside from all this, simply getting an overhead view of the ego vehicle in either case may not result in an accurate BEV ground truth due to the presence of vegetation and other structures (traffic lights, bridges, etc.) that obstruct the overhead view [31].
@@
To overcome these challenges, our paper makes two main contributions. First, we introduce SimBEV, an extensively configurable and scalable randomized synthetic data generation tool that incorporates information from multiple sources to capture accurate BEV ground truth data. In addition to diverse driving scenarios, SimBEV also provides 3D object bounding boxes and supports a variety of configurable perception and localization sensors. This makes it an invaluable tool for empowering researchers to explore various computer vision tasks, and ultimately, accelerate the development of more capable and resilient autonomous driving systems. Second, we use SimBEV to create the SimBEV dataset, a comprehensive large-scale dataset that can serve as a benchmark for a variety of perception tasks and have applications beyond it. One data sample generated by SimBEV is shown in Fig. 1.
@@ 2 Related Work
Real-world driving datasets generally target specific subsets of perception tasks, as the high costs associated with data collection and labeling limit their scope. One of the most prominent and oldest driving datasets is the KITTI dataset [12], which supports depth estimation and 2D and 3D object detection and tracking. Other notable image-based datasets include Cityscapes [6], and Mapillary [44], which are geared towards segmentation, while A*3D [46] focuses on 3D object detection. More recently, large-scale datasets such as BDD100K [71], Waymo Open [54], ApolloScape [18], Argoverse 2 [63], and nuScenes [2] have emerged, offering multi-modal data and multi-task annotations but primarily emphasizing object detection and tracking.
Synthetic driving datasets are compiled using graphics engines and physics simulators. For instance, SYNTHIA [49] incorporates RGB and semantically segmented images generated by its dedicated simulator. Video games have also served as a source of data. For example, GTA-V [48] offers RGB and semantically segmented images extracted from GTA. ViPER [51] expands upon it by including optical flow images and discrete environmental labels. The introduction of CARLA (CAR Learning to Act) Simulator [9], built upon Unreal Engine 4 (UE4), enabled the systemic generation of driving datasets. All-in-One Drive [62] is one such dataset, featuring support for multiple sensors and a focus on SPAD (Single-Photon Avalanche Detector)-lidars. Another notable dataset is the SHIFT dataset [55], a large-scale multi-task multi-modal synthetic dataset for autonomous driving, designed to simulate discrete and continuous shifts in weather and traffic conditions.
Existing datasets offer limited support for BEV segmentation. In nuScenes [2] BEV segmentation is only supported for static map elements (drivable area, lane line, pedestrian crossing, etc.). In Lyft Level 5 [16] and Argoverse 1 and 2 [4, 63], BEV ground truth is obtained through combining map elements and vehicle bounding boxes observable by the perception sensors, limiting BEV ground truth area and potentially missing some occluded objects.
Some vehicle-to-everything (V2X) datasets have support for BEV segmentation, but it is again limited. H-V2X [31] captures the BEV ground truth using overhead cameras installed along a 100 km highway, with the data limited to highway driving and mostly suitable for highway motion forecasting. The synthetic OPV2V [68] dataset provides BEV ground truth only for the drivable area, lane line, and vehicle classes, and the former can be inaccurate because it is calculated using waypoints only. Finally, the synthetic V2X-Sim [24] dataset captures the BEV ground truth using an overhead camera, limiting driving scenarios and causing inaccuracies due to vegetation, traffic light poles, and other structures that obstruct the overhead view.
@@ 3 SimBEV
SimBEV relies on a customized version of CARLA 0.9.15 [9] (see the Supplementary Material) to simulate the environment, perception sensors, and traffic behavior, although it is compatible with the standard release of CARLA as well. In the following, we describe how SimBEV operates and generates driving data.
@@ 3.1 Design
Instead of using carefully crafted scenarios, designing which can be very time-consuming, SimBEV relies on randomizing as many simulation parameters as possible to quickly create a statistically diverse set of scenarios. To create a new dataset, SimBEV generates and collects data from consecutive episodes, or scenes.
To start, the user configures the desired number of scenes for each map (i.e. driving environment, can be an existing CARLA map or a custom one) for each of the training, validation, and test sets. SimBEV checks to see if a SimBEV dataset already exists. If it does, SimBEV subtracts the number of existing scenes in that dataset from the number of desired scenes for each map. This feature allows the user to expand an already existing SimBEV dataset, or to continue with dataset creation in the event of a crash. SimBEV also allows the user to replace individual scenes.
At the start of each scene, SimBEV creates uniformly distributed waypoints within a certain distance of each other (specified by the user, 24.0 m by default) across the map’s roads and selects one at random to use to spawn the ego vehicle and the desired perception and localization sensors attached to it, though the user also has the option to specify a set of spawn points that SimBEV can choose from. SimBEV then configures the weather at random and, if at night, changes the intensity of street lights randomly.
To create background traffic, SimBEV randomly chooses the desired number of background vehicles and pedestrians, although the user has the option to specify each. It then uses the waypoints mentioned previously to spawn random vehicles at random positions, and spawns pedestrians at random on walkable areas of the map. CARLA’s traffic manager controls the behavior of vehicles and pedestrians throughout the episode.
Because all vehicles and pedestrians start from rest, SimBEV runs the simulation for a user-specified period of time (called warm-up duration, by default 4 s) before collecting data for a user-specified period of time (by default 16 s). During data collection, at each time step SimBEV collects and saves data from the desired sensors, and calculates and saves both 3D object bounding boxes and BEV ground truth masks. Following data collection, it saves meta-information about the data and a log of the scene, destroys the vehicles, pedestrians, and sensors, and starts a new scene.
@@ 3.2 Sensors
SimBEV supports a variety of sensors available in CARLA. They include five different camera types (RGB, semantic segmentation, instance segmentation, depth, and optical flow), two different lidar types (regular and semantic), radar, GNSS, and IMU, as shown in Fig. 1. The user has full control over each sensor’s parameters (e.g. a camera’s resolution or FoV), but the placement of the sensors is fixed for now. The cameras are placed at six locations above the vehicle similar to [2] to offer a 360-degree view of the vehicle’s surroundings, while a radar is placed on each side of the vehicle. GNSS and IMU sensors are placed at the center of the vehicle (the origin of the vehicle’s coordinate system), while the lidar is placed high above that center.
@@ 3.3 Scene configuration
As discussed in what follows, numerous parameters are randomized for each scene to ensure that scenes are as unique and diverse as possible.
@@
Weather in CARLA is configured using several parameters such as rain intensity, fog density, sun altitude angle, wind intensity, etc. By default, SimBEV randomly assigns these parameters for each scene to obtain diverse weather conditions, but the user has the option to set any of these parameters to a fixed value. For instance, setting the sun altitude angle to anything less than zero creates night-time scenes.
@@
SimBEV randomly selects background vehicles from CARLA’s vehicle library, which includes sedans, vans, trucks, heavy goods vehicles (HGVs), buses, bicycles, motorcycles, and emergency vehicles (firetruck, ambulance, and police car). Emergency lights are randomly turned on for the latter vehicles, and the user can set the probability that this happens (by default 0.5). When possible, vehicle colors are randomly selected from a set of recommended colors, so e.g. a sedan can have one of 160 standard colors, but the firetruck will always be red. Additionally, some CARLA vehicles support articulated doors, so when these vehicles come to a stop (e.g. at a traffic light), SimBEV may randomly open one or all of their doors.
SimBEV randomly selects each vehicle’s maximum speed (relative to the speed limit, e.g. 10% over or under) and how close that vehicle can get to the vehicle in front of it when stopped. It also randomly selects how long each traffic light stays green. However, the user has the option to set any of these parameters to a fixed value.
SimBEV randomly selects pedestrians from CARLA’s walker library (which contains models of different ages, genders, races, and body types), sets their walking speed at random, and gives them random destinations to go to.
@@
SimBEV allows the user to turn off street and/or building lights (which are automatically turned on at night). Moreover, the user can choose to randomize building light colors, and/or change the intensity of all street lights by a fixed, if desired random, value. Moreover, based on the probability set by the user, SimBEV randomly turns off individual street lights to simulate broken street lights in the real world.
@@
If desired by the user, some vehicles (including the ego vehicle) may drive recklessly, ignoring traffic lights, traffic signs, and collision with other vehicles and pedestrians, as shown in Fig. 2. The user can control the likelihood of reckless driving which can create interesting edge cases. Additionally, the user has control over the share of pedestrians who are allowed to (but may or may not) cross the road at any point, not just at crosswalks.
Figure 2: Reckless ego vehicle running over a cyclist.
3.4 Data annotations
SimBEV currently offers two main annotation types: 3D object bounding boxes and BEV ground truth. The output of some perception sensors (e.g. segmentation, depth, and optical flow cameras) can be used as annotation as well, but they are not discussed here.

### v2

@@ Abstract
Bird’s-eye view (BEV) perception has garnered significant attention in autonomous driving in recent years, in part because BEV representation facilitates multi-modal sensor fusion. BEV representation enables a variety of perception tasks including BEV segmentation, a concise view of the environment useful for planning a vehicle’s trajectory. However, this representation is not fully supported by existing datasets, and creation of new datasets for this purpose can be a time-consuming endeavor. To address this challenge, we introduce SimBEV. SimBEV is a randomized synthetic data generation tool that is extensively configurable and scalable, supports a wide array of sensors, incorporates information from multiple sources to capture accurate BEV ground truth, and enables a variety of perception tasks including BEV segmentation and 3D object detection. SimBEV is used to create the SimBEV dataset, a large collection of annotated perception data from diverse driving scenarios. SimBEV and the SimBEV dataset are open and available to the public.
@@ 1 Introduction
Autonomous driving promises a future with safer, cleaner, more efficient and reliable transportation systems [11, 54]. As development of autonomous vehicle (AV) technology has accelerated in recent years, so has the need for perception algorithms capable of understanding complex driving scenarios in diverse environments [69, 8]. High-quality driving datasets have been at the center of recent progress, serving as a foundation for training and benchmarking novel perception algorithms. It is vital for such datasets to encompass a wide variety of driving scenarios and encapsulate a diverse set of road types, weather conditions, and traffic patterns, so perception models can effectively generalize to real-world situations [31, 49, 15].
@@
As essential in this context is multimodal sensor fusion, which enhances the performance of perception algorithms by compensating for the weaknesses of one modality with the strengths of others [72, 6, 63]. Sensor fusion improves an AV’s understanding of its environment [55] (especially in adverse weather conditions [2]), enables robust decision making in dynamic scenarios [19, 47], and opens the door to perception models capable of performing multiple tasks simultaneously [33, 39, 44, 20]. Consequently, it is imperative for driving datasets to support a wide array of sensors and perception tasks to facilitate the development of multifaceted perception systems that take advantage of the strengths of different sensing modalities.
@@
Bird’s-eye view (BEV) perception has attracted significant attention in recent years for two main reasons [36]. First, BEV representation is conducive to the fusion of information from different modalities, perspectives, and agents, and extracted BEV features can be used for various perception tasks. Second, BEV segmentation offers a concise, geometrically accurate, and semantically rich view of the environment, and can be used by motion planning, behavior prediction, and control algorithms. These two factors have led to the proliferation of perception algorithms that use BEV representation for 3D object detection, BEV segmentation, or both [33, 20, 57, 62, 29, 21, 23, 27, 68, 61, 37, 53, 66, 4, 32, 38, 14, 34, 17, 56, 28, 35, 26, 71, 10, 13, 41, 60, 65, 25, 70].
@@
Figure 1: A data sample generated by SimBEV. The left half depicts a 360-degree view of the ego vehicle’s surroundings using different camera types (from top to bottom RGB, sematic segmentation, instance segmentation, depth, and optical flow cameras, respectively). On the right half, views of lidar, semantic lidar, radar, and the BEV ground truth are shown from top to bottom, respectively. Some images also contain 3D object bounding boxes colored according to the object’s class.
@@
Despite growing interest in BEV perception, few existing datasets support BEV segmentation. For the ones that do, either BEV ground truth is limited to static map elements (drivable area, pedestrian crossing, etc.) [3], or BEV ground truth is only provided for a small window around the ego vehicle (obtained by combining map elements with 3D object bounding boxes [16]) as objects further away may be occluded from the ego vehicle’s view.
@@
Creating a new dataset to fill this gap is a challenging endeavor. Real-world driving data require (at least in part) labor-intensive hand annotation and need to be collected over a long period of time to ensure that weather conditions and traffic patterns present in the dataset are diverse [52, 31]. On the other hand, synthetic driving data often consist of user-designed scenarios that in most cases do not capture the full diversity of the environment. Moreover, simply capturing the overhead view of the ego vehicle in either case may not be enough to obtain the BEV ground truth due to the presence of vegetation and other structures (traffic lights, bridges, etc.) that obstruct that view [30].
@@
To overcome these challenges, our paper makes two main contributions. First, we introduce SimBEV, a synthetic data generation tool based on CARLA Simulator [9] that uses domain randomization to create diverse driving scenarios. SimBEV supports a comprehensive array of sensors and incorporates information from multiple sources to capture accurate BEV ground truth and 3D object bounding boxes. It enables a variety of perception tasks, including BEV segmentation and 3D object detection, making it an invaluable tool for computer vision researchers and helping accelerate the development of more capable autonomous driving systems. Second, we use SimBEV to create the SimBEV dataset, a comprehensive large-scale dataset that can serve as a benchmark for a variety of perception tasks. A data sample generated by SimBEV is shown in Fig. 1.
@@ 2 Related Work
Real-world driving datasets often target specific subsets of perception tasks, as the high costs associated with data collection and labeling limit their scope. One of the oldest and most prominent driving datasets is the KITTI dataset [12], which supports depth estimation and 2D/3D object detection and tracking. Other notable image-based datasets include Cityscapes [7], and Mapillary [40], which are geared towards segmentation, while A*3D [42] focuses on 3D object detection. More recently, large-scale datasets such as BDD100K [67], Waymo Open [50], ApolloScape [18], Argoverse 2 [59], and nuScenes [3] have emerged, offering multi-modal data and multi-task annotations but primarily emphasizing object detection and tracking.
Synthetic driving datasets are compiled using graphics engines and physics simulators. For example, SYNTHIA [46] incorporates RGB and semantically segmented images generated by its dedicated simulator. Video games have also served as a source of data. For instance, GTA-V [45] offers RGB and semantically segmented images extracted from GTA. ViPER [48] expands on GTA-V by including optical flow images and discrete environmental labels. The introduction of CARLA [9] fostered systemic generation of driving datasets. All-in-One Drive [58] is one such dataset, providing support for multiple sensors with a focus on simulating SPAD (Single-Photon Avalanche Detector)-lidars. Another is SHIFT [51], a large-scale multi-task multi-modal dataset for autonomous driving, designed to simulate discrete and continuous changes in weather and traffic conditions to evaluate domain adaptation strategies.
Existing datasets offer limited support for BEV segmentation. In nuScenes [3], BEV segmentation is only supported for static map elements (drivable area, lane line, pedestrian crossing, etc.). In Lyft Level 5 [16] and Argoverse [5, 59], BEV ground truth is obtained by combining map elements and vehicle bounding boxes observable by the ego vehicle’s perception sensors, limiting BEV ground truth area and/or missing occluded objects.
Some vehicle-to-everything (V2X) datasets provide limited support for BEV segmentation as well. H-V2X [30] captures the BEV ground truth using overhead cameras installed along a 100 km highway, with data limited to highway driving and mostly suitable for highway motion forecasting. CARLA-based OPV2V [64] dataset provides the BEV ground truth, but only for the drivable area, lane line, and vehicle classes, and the one for drivable area can be inaccurate due to its sole reliance on CARLA-generated waypoints. Finally, CARLA-based V2X-Sim [24] dataset captures the BEV ground truth for several classes using an overhead camera, which can be inaccurate due to the presence of vegetation, traffic light poles, and other structures that obstruct the overhead view.
@@ 3 SimBEV
Figure 2: SimBEV’s logic flow when creating a new dataset. The arrow exiting green nodes at the top indicates the action taken when the condition in that node is no longer satisfied.
SimBEV relies on CARLA 0.9.15 [9] equipped with a custom content library (see the Supplementary Material) to simulate the environment, perception sensors, and traffic behavior, although it is compatible with the standard release of CARLA as well. SimBEV streamlines, automates, and manages the entire data collection process for the user (who controls it through a single configuration file) by manipulating simulation elements through CARLA’s Python API. It equips the user with CARLA’s customizability when necessary and takes charge when not. This flexibility and ease of use enables researchers to quickly create custom datasets that suit their needs.
@@ 3.1 Design
SimBEV works by randomizing (within some bounds) as many simulation parameters as possible to create a statistically diverse yet realistic set of scenarios. To create a dataset, SimBEV generates and collects data from consecutive episodes, or scenes, each with a unique configuration.
SimBEV’s logic flow is shown in Fig. 2. To start, the user configures the desired number of scenes for each map (i.e., the driving environment, can be an existing CARLA map or a custom one) for the training, validation, and test sets. SimBEV checks to see if a SimBEV dataset already exists. If so, it subtracts the number of existing scenes in that dataset from the number of desired scenes for each map. This lets the user expand an already existing SimBEV dataset, or continue with dataset creation in the event of a crash. SimBEV also lets the user replace individual scenes.
At the start of each scene, SimBEV creates uniformly distributed waypoints a certain distance (specified by the user) apart from each other across the map’s roads. It selects one at random and spawns the ego vehicle and attached sensors there, though the user has the option to specify a set of spawn coordinates that SimBEV must choose from instead. SimBEV then configures the weather randomly and, if at night, changes the intensity of street lights at random.
For background traffic, SimBEV randomly selects the desired number of vehicles and pedestrians, although the user can specify each. SimBEV then uses the waypoints mentioned earlier to spawn random vehicles at random locations, and spawns pedestrians randomly on walkable areas of the map. CARLA’s traffic manager controls the behavior of vehicles and pedestrians throughout the scene.
Because all vehicles and pedestrians start from rest, SimBEV runs the simulation for a user-specified period of time (called warm-up duration) to reach a more realistic state, before collecting data for a user-specified period of time. During that period, SimBEV saves data from the desired sensors at each time step, and calculates and saves both 3D object bounding boxes and the BEV ground truth. Following that, it saves meta-information about the collected data and a log of the scene, destroys the vehicles, pedestrians, and sensors, and moves on to the next scene.
@@ 3.2 Sensors
SimBEV supports a variety of sensors available in CARLA, including five different camera types (RGB, semantic segmentation, instance segmentation, depth, and optical flow), two different lidar types (regular and semantic), radar, GNSS, and IMU, as shown in Fig. 1. The user has full control over each sensor type’s parameters (e.g. a camera’s resolution or FoV), but the placement of sensors is fixed for now. Similar to [3], cameras are placed at six locations above the vehicle to offer a 360-degree view of the vehicle’s surroundings, while a radar is placed on each of the four sides of the vehicle. The GNSS and IMU are placed at the center of the vehicle (the origin of the vehicle’s coordinate system), and a lidar is placed high above that center.
@@ 3.3 Scene configuration
As will be discussed in what follows, numerous parameters are randomized for each scene to ensure that generated scenes are as unique and diverse as possible.
@@
Weather in CARLA is controlled using several parameters such as fog density, sun altitude angle, wind intensity, etc. By default, SimBEV randomly selects these parameters for each scene to create diverse weather conditions (subject to some constraints to ensure the realism of the weather), but the user also has the option to set any of the parameters to a fixed value. For instance, setting the sun altitude angle to anything less than zero creates a dataset of night-time scenes.
Figure 3: In a scene generated by SimBEV, a reckless ego vehicle runs over a cyclist.
@@
SimBEV randomly selects background vehicles from CARLA’s vehicle library, which includes sedans, vans, trucks, heavy goods vehicles (HGVs), buses, bicycles, motorcycles, and emergency vehicles, whose emergency lights are turned on randomly. When possible, vehicle colors are selected at random from a vast set of available colors (e.g. a sedan can change colors but a firetruck will only be red). Some vehicles support articulated doors, so when these vehicles come to a stop - e.g. at a traffic light - SimBEV may randomly open one or all of their doors.
SimBEV randomly chooses each vehicle’s maximum speed (relative to the speed limit, e.g. 10% over/under) and how close vehicles can get to each other when coming to a stop. It also randomly selects how long each traffic light stays green. However, the user always has the option to set any of these parameters to a fixed value.
SimBEV randomly chooses pedestrians from CARLA’s walker library (which contains models of different age, gender, race, and body type), sets their walking speed at random, and gives each a random destination to go to.
@@
SimBEV gives the user the option to turn off all street and/or building lights at night. It also lets the user randomize building light colors, and/or change the intensity of all street lights by a fixed, if desired random, value. In addition, SimBEV randomly turns off individual street lights based on a probability set by the user to simulate broken street lights in the real world.
@@
If desired by the user, some vehicles (including the ego vehicle) can drive recklessly, ignoring traffic lights, traffic signs, and collisions with other vehicles and pedestrians, as shown in Fig. 3. The user controls the likelihood of reckless driving, which can result in interesting edge cases. The user also has control over the share of pedestrians allowed to jaywalk and cross the road at any point, not just at crosswalks.
Figure 4: Ground elements (roads, sidewalks, etc.) in CARLA use one-way visible materials, appearing invisible to a camera placed below them. We use this property to capture accurate BEV ground truth by placing a camera below the ego vehicle looking up.
3.4 Data annotation
SimBEV offers two main types of data annotation: 3D object bounding boxes and BEV ground truth. The output of some perception sensors such as segmentation, depth, and optical flow cameras and semantic lidar can serve as annotation as well, but we do not discussed them here.